OpenStudy (anonymous):

The stiffness of a beam of rectangular cross section is proportional to the breadth and the cube of the depth. If the breadth is 20 cm, express the stiffness as a function of the depth.

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

there is no choices

jhonyy9 (jhonyy9):

i am not sure that i understand it right but my opinion is let s = sriffness of a beam b = breadth cube of the depths = d so we know that breadth = 20 cm and that s is proportional to the b and d so than now can you writing this proportionalitie of these terms ?

jhonyy9 (jhonyy9):

are you here ?

jhonyy9 (jhonyy9):

the cube of the depth mean the depth cubed so d^3 ?

jhonyy9 (jhonyy9):

so than we can wriiting s/b = s/d^3 s/20 = s/d^3 s^2 = 20d^3 s = sqrt(20d^3) = 2dsqrt(5d) s= 2dsqrt(5d) sorry but i am not sure that i have understood it all right